OpenStudy (anonymous):

how do you do (-2,4) (3,-6) If there is no slope. Do i have to find the slope first and than write the equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so slope i got is m= -2/-8 what now

OpenStudy (anonymous):

use the point slope form

OpenStudy (anonymous):

which is

OpenStudy (anonymous):

I am stuck on y= -2/-8x +

OpenStudy (anonymous):

started by y-y1=m (x-x1) y-4=-2/-8 (x+2)

OpenStudy (anonymous):

right, now notice that you can reduce the slope to 1/4

OpenStudy (anonymous):

then just distribute the 1/4 into the parentheses

OpenStudy (anonymous):

can u show me

OpenStudy (anonymous):